Katy:

Circle O has a diameter of 15 cm. Points T and P lie on the circle, and mPT=180. Find the area of sector TOP. Leave your answer in terms of π. 7.5π square centimeters 15π square centimeters 28.125π square centimeters 112.5π square centimeters

Razor:

Well the area of a sector formula is: A = (degree/360) × πr² We know the degree so keep that in mind. The diameter is 15  the radius is half of the diameter so the radius is 7.5. Now we plug these into the equation: A = (180/360) × π7.5² A= (1/2) × π56.25 A= 28.125 × π A = 28.125π